A gas has a volume of 3.0 L at 25°C what will the volume of this gas in L if the temperature is increased to 79°C. Assume pressure is constant.


1
Expert's answer
2021-03-29T05:15:36-0400

pV=nRT, n= pV/RT = 3 x 1 /8.314 x 298 = 0,00121086289.

V= nRT/p = 0,00121086289 x 8.314 x 352/1 = 3,54362416 L.